media%2F66c%2F66cb1845-8ec9-457e-a591-48
Answer #1

part a:

for string 1, wave velocity=sqrt(tension/mass per unit length)

=sqrt(20*9.8/0.005)=198 m/s

part b:

fundamental frequency of first string=wave speed/(2*length )

=198/(2*0.5)=198 Hz

first three harmonics are:

1st: 198 Hz

2nd: 198*2=396 Hz

3rd : 594 Hz

part c:

sketch:

fundamental 0.5 0.0005 0.001 econic 0.0025 0.003 0.5 0.0005 0.001 0.0015 0.002 0.0025 0.003 0.5 0.0005 0.001 0.0015 0.002 0.0

part d:

tension in the string=weight of the block=M*g

speed of wave=sqrt(tension /mass per unit length)

=sqrt(M*g/0.005)

third harmonic frequency=3*speed/(2*length)

=3*sqrt(M*g/0.005)/(2*0.5)

=3*sqrt(M*g/0.005)

part e:

as beat frequency is 2 per second, frequency in second string is either 198+2=200 Hz or 198-2=196 Hz

then 3*sqrt(M*g/0.005)=200

==>M=2.2675 kg

or 3*sqrt(M*g/0.005)=196

==>M=2.178 kg

so two possible values of M can be 2.2675 kg or 2.178 kg.